Annuity( Stept)

I am looking not for the answer. but more so the steps. My textbook gave the answer of 104.223.28. I know one of my steps is wrong and I am getting a different answer. Payment 3500, PayPer-semiannually, Intr 8% Compsemiannually,length 10yrs- what is the FV?
Answer #1

I figured it out: (1+0.04)20-1= (1.04)20/0.04 =2.1911-1/0.04= 1.1911/0.04= 29.778

29.778*3500=104.221.25 close enough

source: Business Mathematics for College
answered by: anonymous
Answer #2

SOLUTION :


r per semi-annual period = 8/2 = 4% = 0.04

=> (1 + r) = 1.04

n = 10*2 = 20 semi-annual periods.

A = 3500 (assumed as $)


So,


FV 

=  A((1 + r)^n - 1) /  r 

= 3500(1.04^20 - 1) / (0.04)

= 104223.28  ($)  (ANSWER).

answered by: Tulsiram Garg

> Last calculation in the question part of the student : 3500*29.778 is not correct. It should be 104223.78 . Student may recheck his calculation.

Tulsiram Garg Wed, Dec 8, 2021 11:27 PM